# 6D Rigid Body Localization and Velocity Estimation via Gaussian Belief Propagation

### v2

@@ V-B Convergence Behavior and Complexity Analysis
Prop. Stationary Linear GaBP (Alg. 1) $O(NMK)$ $0.15$ms
Prop. Moving Linear GaBP (Alg. 3) $O(NMK)$ $0.14$ms
SotA Position/Velocity Est. [30] $O(MNK^{2})$ $1.6$ms
Prop. Stationary Double GaBP (Alg. 2) $O(NMK^{2})$ $1.1$ms
SotA Stationary Parameter Est. [11] $O(K^{3}+MN)$ $0.24$ms
Prop. Moving Double GaBP (Alg. 4) $O(NMK^{2})$ $1$ms
SotA Moving Parameter Est. [11] $O(K^{3}+MN)$ $0.39$ms
First, Figure 9 illustrates the convergence behaviour of the four proposed algorithms, where it can be observed that all four algorithms converge very efficiently, with the bivariate Algorithm 1 and 3 converging slightly faster and to a higher estimation accuracy than the two linear algorithms. It should also be noted that since the SotA algorithms are not iterative, a convergence analysis comparison is not relevant.
Finally, the computational complexities of the proposed algorithms have been outlined in Table I in terms of the complexity order on the system size parameters using the well-known Big-O notation, as well as a convenient measure of practical runtimes in seconds simulated via an ordinary computer.
@@ VI Conclusion
We presented a novel and efficient framework for solving the sixth-dimensional (6D) RBL problem, for under stationary and moving i.e., estimation of 3D location and 3D rotation, as well as corresponding translation and angular velocities, via a series of tailored GaBP message passing estimators. First, a linear GaBP for the sensor position estimation based on the obtained range estimates from the wireless sensors is derived, from which the RBL system is reformulated via the small-angle approximation to enable the construction of a second bivariate GaBP which is capable of directly estimating the rotation angles and the translation distances. Next, Doppler measurements were obtained, designing a system of linear equations that can be used to estimate the sensor velocity via a linear GaBP, followed by the estimation of angular and translational velocity by the construction of another bivariate GaBP. The proposed RBL method is shown to outperform the SotA method in all of the position, rotation, translation, sensor velocity, angular velocity and translational velocity estimation performance, in addition to the complexity advantage highlighted by the analysis.

### v3

@@ V-A RBL Parameter Estimation Performance Against the SotA
Prop. Stationary Linear GaBP (Alg. 1) $O(NMK)$ 0.150.15ms
Prop. Moving Linear GaBP (Alg. 3) $O(NMK)$ 0.140.14ms
SotA Position/Velocity Est. [32] $O(MNK^{2})$ 1.61.6ms
Prop. Stationary Double GaBP (Alg. 2) $O(NMK^{2})$ 1.11.1ms
2-Stage SotA Stationary Parameter Est. [12] $O(K^{3}+MN)$ 0.240.24ms
SDR-based SotA Stationary Parameter Est. [30] $O\left((K+1)\left(MN(K+1)^{4}+M^{2}N^{2}(K+1)^{2}+(K+1)^{8}\right)\right)$ 4.64.6ms
Prop. Moving Double GaBP (Alg. 4) $O(NMK^{2})$ 11ms
2-Stage SotA Moving Parameter Est. [12] $O(K^{3}+MN)$ 0.390.39ms
SDR-based SotA Moving Parameter Est. [30] $O\left(8\sqrt{2}(K+1)\left(MN(K+1)^{4}+M^{2}N^{2}(K+1)^{2}+4(K+1)^{8}\right)\right)$ 12.212.2ms
Next, the second set of results evaluates the performance of the proposed rigid body motion estimation algorithms within the moving rigid body scenario. First, Figure 9 illustrates the RMSE of the sensor velocity estimates of Algorithm 3 compared to the SotA solution of [32]. It can be observed that the proposed method outperforms the SotA in all velocity errors with a large performance gain. In turn, Figures 9 and 10 show the performance of the estimate velocity parameters, $i.e.$, the translational velocity and the angular velocity.
Figure 11: Convergence of the proposed algorithms for varying noise levels of $\sigma=10^{0}$ and $\sigma=10^{-2}$.
In Figure 9, the translational velocity estimate of the proposed method in Algorithm 4 is compared to the SotA methods of [12] and [30]. It can be observed that as previously, the proposed method outperforms the SotA for all range errors. In turn, Figure 10 illustrates the performance of the angular velocity estimation, where similar to the stationary rotation estimation results, the proposed method outperforms the 2 Stage SotA of [12] for all range errors, with a larger gain reaching the MFB for high noise and a smaller gain for low noise, while the SDR-based method [30] outperforms our proposed method for low noise levels.
V-B Convergence Behavior and Complexity Analysis
Here, the performance of the proposed RBL parameter estimation algorithms are further assessed in terms of the computational complexity and convergence behavior, which is thoroughly investigated in [50, 51, 52], and compared against the SotA methods. First, Figure 11 illustrates the convergence behavior of the four proposed algorithms, where it can be observed that all four algorithms converge very efficiently, with the bivariate Algorithms 2 and 4 converging slightly steeper, and to a higher estimation accuracy than the two linear algorithms. It should also be noted that since the SotA algorithms are not iterative, a convergence analysis comparison is not relevant.
Finally, the computational complexities of the proposed algorithms have been outlined in Table I in terms of the complexity order on the system size parameters using the well-known Big-O notation, as well as a convenient measure of practical runtimes in seconds simulated via an ordinary computer. It can be observed that the proposed GaBP algorithms provide an accurate estimation of the RBL parameters with a low computational complexity, comparable to the SotA methods of [12] and [32], while being far superior to the SotA method of [30].
@@ VI Conclusion
We presented a novel and efficient framework for solving the sixth-dimensional (6D) RBL problem, for under stationary and moving i.e., estimation of 3D location and 3D rotation, as well as corresponding translation and angular velocities, via a series of tailored GaBP message passing estimators. First, a linear GaBP for the sensor position estimation based on the obtained range estimates from the wireless sensors is derived, from which the RBL system is reformulated via the small-angle approximation to enable the construction of a second bivariate GaBP which is capable of directly estimating the rotation angles and the translation distances. Next, Doppler measurements were obtained, designing a system of linear equations that can be used to estimate the sensor velocity via a linear GaBP, followed by the estimation of angular and translational velocity by the construction of another bivariate GaBP. The proposed RBL method is shown to outperform the SotA method in all the position, rotation, translation, sensor velocity, angular velocity and translational velocity estimation performance, in addition to the complexity advantage highlighted by the analysis.